OpenStudy (anonymous):

Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?

OpenStudy (anonymous):

let dimes = x, quarters =100 -x x*10 + (100-x) * 25 = 1975 => 2500 - 15x = 1975 => x = 525/ 15 = 35 so, dimes 35, quarters 65